Winston bakes 4 pies per hour, and Clyde bakes 6 pies per hour. If Winston bakes alone for one hour, and then Clyde joins him, how many minutes will it take them to bake 32 pies?
Answer by nerdybill(7384)

Winston bakes 4 pies per hour, and Clyde bakes 6 pies per hour. If Winston bakes alone for one hour, and then Clyde joins him, how many minutes will it take them to bake 32 pies?
.
Let x = hours it takes to bake 32 pies
then
1(4) + x(4 + 6) = 32
4 + x(10) = 32
4 + 10x = 32
10x = 28
x = 2.8 hours
.
Converting to minutes:
x = 2.8(60)
x = (2 + .8)(60)
x = 2(60) + .8(60)
x = 120 + 48
x = 168 minutes